Question 1 A rail shuttle line with a length L-3250m has 6 stations. The travel time between segments and dwell times at each station are given in the table below, you may assume the running times and dwell times are the same in both directions (A to F and F to A). This shuttle service operates at 6-min headway. The minimum terminal time is 150 sec at each terminus. Dwell time (sec) 20 Station Segment A-B B-C C-D D-E E-F Running time (sec) 120 80 100 120 60 40 40 40 20 (i) Determine the operating speed and cycle speed for this light rail line. (ii) Construct a timetable schedule (in hh:mm:ss) showing departure time at each station for this line covering the period 08:00:00 to 08:30:00. The first train should depart at A 08:00:00
Answer #1

Total time taken from A-F is 120+80+100+120+60+40+60+40+40=660 seconds

A) operating speed = 3250/660 = 4=92m/s

Cycle time = 480+240+150*2 = 1020 seconds

B)arrival time arrival time towarda A Departure time Departure time towards A Segment Running time Station Owell time towarda F towards F A-8 B-C C-D D-E E-F 120 А 80 B 100 C 120D 60 E 20 40 60 40 40 40 7:59:40 8:05:40 8:1110 8:17210 8:23:40 8:02:00 9:08:00 8:14:00 8:20:00 8:26:00 8:04:00 8:10:00 8:16:00 8:22:00 8:28:00 8:06:40 8:12:40 8:18:40 8 24:40 8:30:40 8:00:20 :15:20 8:21:20 8:27:20 :33:20 8:1100 8:17:00 8:23:00 8:29:00 8:35:00 8:00:00 8:02:40 8:05:00 8:07:20 8:10:00 8:06:00 8:12:00 8:18:00 8:24:00 3:09:40 8:14:40 8:20:40 8:26:40 8:11:00 8:17:00 8:23:00 8:29:00 8:13:208:19:20 8:25:20 8:31:20 3:16:00 3:22:00 :2B:00 8:34:00 8:25:10 8:22:30 8:20:10 8:17:50 B:15:10 8:13:30 8:31:10 8:23:30 8:26:10 8-23:50 8:21:10 8:19:30 8:37:10 8:34:30 8:32:10 8:29:50 8:27:10 8:25:30 8:23:10 8:29:10 8:35:10 8:2110 8:27:10 8:33:10 8:18:30 8:24:30 8:30:30 8:15:50 8:21:50 8:27:50 8:14:10 8:20:10 826:10